OpenStudy (anonymous):

Solve for x. 8x/7 + 9 = 30 x = 18 x = 24 x = 34 x = 44

OpenStudy (anonymous):

IF I'M RIGHT MEDALS AND FANS: Solution for 8x/7+9=30 equation: 8x/7+9=30 We simplify the equation to the form, which is simple to understand 8x/7+9=30 Simplifying: +1.14285714286x+9=30 We move all terms containing x to the left and all other terms to the right. +1.14285714286x=+30-9 We simplify left and right side of the equation. +1.14285714286x=+21 We divide both sides of the equation by 1.14285714286 to get x. x=18.375